### v2

# Quantum Reinforcement Learning Trading Agent for Sector Rotation in the Taiwan Stock Market

Abstract
We propose a hybrid quantum-classical reinforcement learning framework for sector rotation in the Taiwan stock market. Our system employs Proximal Policy Optimization (PPO) as the backbone algorithm and integrates both classical architectures (LSTM, Transformer) and quantum-enhanced models (QNN, QRWKV, QASA) as policy and value networks. An automated feature engineering pipeline extracts financial indicators from capital share data to ensure consistent model input across all configurations. Empirical backtesting reveals a key finding: although quantum-enhanced models consistently achieve higher training rewards, they underperform classical models in real-world investment metrics such as cumulative return and Sharpe ratio. This discrepancy highlights a core challenge in applying reinforcement learning to financial domains—namely, the mismatch between proxy reward signals and true investment objectives. Our analysis suggests that current reward designs may incentivize overfitting to short-term volatility rather than optimizing risk-adjusted returns. This issue is compounded by the inherent expressiveness and optimization instability of quantum circuits under Noisy Intermediate-Scale Quantum (NISQ) constraints. We discuss the implications of this reward-performance gap and propose directions for future improvement, including reward shaping, model regularization, and validation-based early stopping. Our work offers a reproducible benchmark and critical insights into the practical challenges of deploying quantum reinforcement learning in real-world finance.
I Introduction
Financial markets are inherently dynamic and noisy [chiarella2011analysis], driven by complex interactions among macroeconomic cycles, sector-specific trends, and investor behaviors. Sector rotation, the practice of reallocating capital between industry sectors based on their expected performance across economic phases, is a prominent strategy for generating alpha and managing risk in volatile markets. Reinforcement learning (RL) [kaelbling1996reinforcement] offers a powerful paradigm for this challenge. Its ability to learn optimal decision policies through direct interaction with an environment aligns naturally with the sequential, stochastic nature of financial decision-making. Unlike traditional rule-based or supervised learning methods that depend on static datasets, RL agents dynamically adapt their strategies to maximize cumulative rewards over time.
Recent advances in quantum machine learning (QML) [biamonte2017quantum] introduce new possibilities for enhancing these learning systems. By integrating quantum principles like superposition and entanglement, quantum reinforcement learning (QRL) [dong2008quantum, chen2022variational] aims to create more expressive policy representations and accelerate optimization. However, the practical application of QRL is currently constrained by the limitations of Noisy Intermediate-Scale Quantum (NISQ) hardware, including significant noise, shallow circuit depth, and barren plateau phenomena [chen2020variational, kuo2021quantum]. In this work, we investigate the application of a hybrid quantum-classical RL framework to the task of sector rotation within the Taiwan stock market. We have developed a flexible agent based on Proximal Policy Optimization (PPO) [schulman2017proximal] that supports both classical network backbones, such as LSTM [hochreiter1997long] and Transformer [vaswani2017attention], and quantum-enhanced backbones, including QNN [mitarai2018quantum], QRWKV [chen2025qrwkv], and QASA [chen2025quantum]. All agents were trained and backtested under consistent experimental conditions using sector-level capital share data and automated feature engineering. A key finding of our study is that while quantum models frequently achieved higher rewards during training, they consistently underperformed classical models on real-world investment metrics like cumulative return and the Sharpe ratio. This highlights a fundamental disconnect between proxy reward functions and actual investment goals in financial RL applications. Furthermore, it raises critical questions about the interplay between quantum expressiveness, training instability, and reward design within the constraints of NISQ-era technology.
Our contributions are threefold. First, we introduce a reproducible QRL benchmark for sector rotation using real financial data. Second, we provide a systematic comparison of classical and quantum-enhanced policy networks within a shared PPO framework. Third, we analyze the observed discrepancy between training rewards and investment performance, proposing actionable improvements for reward shaping and regularization to bridge this gap.
II Related Work
II-A Sector Rotation Strategies in Financial Markets
Sector rotation is a dynamic and adaptive investment strategy that systematically reallocates capital among various industry sectors based on expectations of future economic performance [cavaglia2001risks]. The underlying principle assumes that different sectors respond differently to various phases of the economic cycle. For example, technology and consumer discretionary sectors often outperform during economic expansions, while defensive sectors such as utilities and consumer staples tend to be more resilient during contractions. Conventional approaches to sector rotation can be categorized into top-down and bottom-up methodologies. The top-down approach involves analyzing macroeconomic indicators such as gross domestic product growth, interest rates, and inflation to infer sector performance trends. In contrast, the bottom-up approach focuses on evaluating the fundamentals of individual companies within each sector to identify opportunities. A combined approach seeks to synthesize both macroeconomic and microeconomic perspectives to form a holistic investment strategy. More recently, quantitative techniques have been employed to enhance sector rotation, including factor-based models that exploit patterns such as momentum, mean reversion, and volatility clustering. However, many of these methods remain static or rule-based, lacking the ability to adapt autonomously to new market conditions or structural changes. This motivates the exploration of reinforcement learning as a means of enabling adaptive, data-driven sector allocation strategies.
II-B Reinforcement Learning in Portfolio Management
Reinforcement learning (RL) provides a computational framework in which an agent learns to make sequential decisions through interactions with a dynamic environment. In the context of financial markets, RL has been applied to various tasks such as portfolio rebalancing, asset allocation, algorithmic trading, and hedging strategies. By learning policies that maximize long-term cumulative rewards, RL agents are capable of uncovering investment strategies that adapt over time. Compared to traditional approaches such as the Markowitz mean-variance optimization framework, RL does not rely on strong statistical assumptions about asset return distributions or covariance matrices. This makes RL particularly well-suited for financial environments, which are often characterized by high volatility, structural breaks, and non-stationary behavior. Furthermore, the incorporation of deep neural networks in deep reinforcement learning (DRL) enables the modeling of nonlinear and temporally dependent financial patterns. Notably, architectures such as LSTM networks and Transformer encoders have been used to capture complex dependencies in sequential financial data, improving an agent’s ability to anticipate market dynamics. However, despite these advances, one of the persistent challenges in applying RL to financial decision-making lies in the design of appropriate reward functions. Agents may overfit to proxy rewards that correlate poorly with true investment objectives, such as risk-adjusted returns, resulting in strategies that perform well during training but fail during backtesting or live trading.
II-C Quantum Machine Learning for Financial Optimization
Quantum machine learning (QML) is an emerging field that explores the integration of quantum computation with machine learning algorithms. By harnessing quantum mechanical phenomena such as superposition, entanglement, and quantum parallelism, QML aims to improve learning efficiency, representational power, and solution quality, especially for high-dimensional or computationally intractable problems. In finance, QML has been proposed for a wide range of applications, including portfolio optimization, option pricing, credit risk modeling, and market simulation. Hybrid quantum-classical models, which combine variational quantum circuits with classical neural networks, have become a practical choice for current hardware capabilities, particularly in the Noisy Intermediate-Scale Quantum (NISQ) era. These models seek to leverage quantum expressiveness while maintaining the stability and scalability of classical computation. Quantum reinforcement learning (QRL), a subfield of QML, focuses on embedding quantum components within RL architectures, such as representing policies or value functions using quantum circuits. Several studies have demonstrated the theoretical benefits of QRL in synthetic environments, showing potential improvements in learning capacity and sample efficiency. Nevertheless, empirical evaluations on real-world tasks remain limited. Key challenges include the difficulty of training quantum models due to barren plateau phenomena, susceptibility to noise and decoherence, and the absence of standardized quantum-compatible financial environments. This work contributes to the growing body of literature by presenting a practical and reproducible QRL framework applied to the real-world problem of sector rotation. In contrast to prior research that primarily focuses on proof-of-concept simulations, our study offers a direct comparison between classical and quantum-enhanced models under consistent financial conditions and highlights the mismatch between training reward optimization and actual investment performance.
III Methodology
III-A Problem Formulation
We formulate the sector rotation problem as a sequential decision-making task in which an agent selects an allocation across multiple industry sectors to maximize long-term investment return. At each discrete time step , the agent observes a financial state vector , derived from engineered features of sector-level market data. Based on this state, the agent chooses an action , representing the target portfolio allocation across sectors. After executing the action, the environment transitions to a new state and returns a scalar reward based on the investment outcome. The goal is to learn a policy that maximizes the expected cumulative discounted reward .
III-B Framework Overview
We adopt a hybrid quantum-classical reinforcement learning framework based on the Proximal Policy Optimization (PPO) algorithm. PPO is a widely used policy-gradient method known for its sample efficiency and training stability. It optimizes a clipped surrogate objective to prevent large policy updates and incorporates entropy regularization to encourage exploration. Our framework supports interchangeable policy and value network backbones, allowing both classical and quantum-enhanced models to be evaluated under consistent training protocols.
III-C Policy Network Architectures
III-C1 Classical Baselines
We implement two classical neural network architectures as baselines:
LSTM [hochreiter1997long]: The Long Short-Term Memory (LSTM) network is a type of recurrent neural network (RNN) designed to capture long-term dependencies in sequential data. For an input sequence , the LSTM updates its hidden state and cell state at each time step using the following equations:
where is the sigmoid function, denotes element-wise multiplication, and , , are learnable parameters.
Transformer [vaswani2017attention]: The Transformer model relies entirely on attention mechanisms to model sequential data, without recurrence. Given an input sequence , each position attends to all others via scaled dot-product attention. The attention for a query , key , and value is computed as:
where , , are learned projections and is the dimension of the key vectors. Multi-head attention extends this by running attention heads in parallel:
After attention, each layer applies a position-wise feed-forward network (FFN) and residual connections:
The final output is passed through layer normalization and stacked over multiple layers to form deep representations.
III-C2 Quantum-Enhanced Variants
We implement three quantum-enhanced variants by incorporating variational quantum circuits (VQCs) into the policy and/or value networks: